# Cocoa: Co-Planning and Co-Execution with AI Agents

### v2

@@ 2.3. AI-Powered Tools for Scientific Research
Is the generation of full research artifacts (e.g., research questions, proposals) the most promising path to assist researchers? We question this assumption in our work. Indeed, when working with other (human) collaborators, researchers often find richness in co-evolving partially completed artifacts (Star and Griesemer, 1989; Lee, 2007; Shrum et al., 2007; Jirotka et al., 2013) and answering or asking questions that stimulate critical thinking and reflection (Overholser, 1993; Padesky, 1993). Another reason for the lack of uptake in AI-generated ideas is the lack of human involvement—researchers do not have opportunities to intervene or provide feedback to the fully automated generation pipelines. A limited number of works have thus explored human-AI co-creation of research artifacts (Babaian et al., 2002; Liu et al., 2024a; Nigam et al., 2024; Kang et al., 2022b). These include in-document commands that trigger an AI assistant to complete a partial citation based on the user’s preferred bibliographies and paper collections (Babaian et al., 2002), interactive node-based editors to iteratively expand upon and refine AI-generated research questions (Liu et al., 2024a; Pu et al., 2024), and emulating colleague and mentor personas using LLMs to work with the user in developing research proposals (Nigam et al., 2024).
@@
However, despite these systems, there is still little clarity on how researchers want to co-create with AI during research planning and execution. The discussed works make assumptions on this front—for example, Liu et al. (Liu et al., 2024a) assumed that researchers require AI assistance with research question formulation, while Nigam et al. (Nigam et al., 2024) assumed that generating a full proposal consisting of a research problem, methods, and experiments would be most helpful to researchers. Rather than focusing on generating one type of research artifact, we contribute an interaction design pattern that allows for flexible specification of the final artifact and collaboration between a researcher and an AI agent that can also yield fruitful intermediate artifacts along the way.
@@ 3. Formative Study
Given the opportunities surfaced from prior work, we aim to improve collaboration between humans and AI agents by leveraging plans as a shared operational representation. We focus on planning for scientific research as a use case. Scientific research presents a challenging planning scenario for both humans and agents due to its complex, multi-step, and fluid workflows and the nuanced decisions researchers make upon processing large amounts of information from information-dense scientific papers. Specifically, we use project documents—a central, running document in which researchers may jot down project ideas, updates, to-dos, meeting notes, and more—as the agent environment. We thus conducted a formative study to answer the following research questions:
  • RQ1:
    What are the properties and opportunities of researchers’ project documents?
  • RQ2:
@@
    • RQ2.1:
      How can we characterize researchers’ intentions for planning?
    • RQ2.2:
      How can we characterize the steps researchers include in their plan to meet those intentions?
  • RQ3:
    How would researchers prefer to interact with an AI-powered co-planning agent in their project documents?
3.1. Participants and Procedure
We recruited 9 Ph.D. students (detailed demographic in Table 1 of Appendix B) through an interest form sent to a research organization’s internal Slack channel and the authors’ professional connections. We targeted Ph.D. students and postdoctoral researchers as they are often the primary owners of research project documents and lead the detailed planning and execution of research projects.
All studies were conducted virtually over Google Meet and lasted around 60 minutes. Prior to the study, we collected, from each participant, an active or past real-world research project document and a brief description of their research interests. The study was broken down into 3 parts: an activity involving the project document shared with us to better understand their current planning behavior (25 minutes), an activity with a Wizard-of-Oz (WoZ) design probe to explore how researchers generate project proposals with new ideas in a document interface with LLM support (25 minutes), and a concluding exit interview where researchers reflected on their experience with the probe and using LLMs in research more generally (10 minutes). Each participant was given a \$35 USD honorarium after the study. The study was reviewed and approved by our organization’s internal IRB. Further procedural details can be found in Appendix C.
3.2. Data Analysis
To answer RQ1, the first author used a hybrid inductive-deductive coding process (Fereday and Muir-Cochrane, 2006) to code participants’ submitted project documents. The first 5 documents were inductively coded to surface common structural elements before the elements were deductively applied to the remaining 4. We iterated on existing elements and added new ones as needed. For RQs 2.1 and 2.2, we performed inductive thematic analysis on two documents—participants’ submitted project documents and in-study planning documents. For RQ2.1, the first author sourced intentions for planning within submitted project documents by locating the snippet of text that initiates a plan. For example, if the plan consists of a bulleted list, the intention is often expressed in the lead-in text that immediately precedes the list. We note that an individual plan item can also signal planning intent if it contains a nested plan. The first author performed thematic analysis by inductively coding the extracted text. For RQ2.2, the first author identified and extracted plan steps participants wrote in both documents before inductively coding them. For RQ3, the first author performed open coding on the study transcripts before thematically analyzing the coded snippets. To enrich the data, the first author also extracted and inductively coded all requests to the assistant from participants’ planning documents. The codes were discussed and iterated upon with other project team members on a weekly basis.
@@ 3.3. Findings
We present the results of our thematic analysis across study transcripts, project documents, and planning documents. We redact any project-specific details for privacy and intellectual confidentiality.
3.3.1. [RQ1] Project documents were continually updated for planning and progress tracking
We examined participants’ real-world research project documents and found several interesting characteristics. First, they were a running planning document that is continually updated throughout a project’s lifecycle. Some documents (P1, P4–6) were in a diary-like format with explicit dates, and others also had clear indicators of chronology, such as having “Current” and “Archive” sections. The content of these documents commonly included meeting notes, planned to-do items, progress updates tied to to-do items, research questions to discuss with collaborators, and links to other documents detailing a particular project component in further depth.
Plans in their documents also had varying degrees of specificity and completion. Some had steps that were concretely listed and culminated in a result at the end of the plan. For example, P2 listed out concrete experimental steps, followed by figures with bar charts showing experimental results. Other plans were partially specified and in-progress. For example, P1’s document contained many partially specified plans with steps such as “what if we focus on only [concept]?” and “maybe [approach] possible?” Finally, documents sometimes served as “scratch paper” for reasoning and high-level goals without concrete plan steps. Most documents (P1, P2–4, P7–P9) contain one or more problem statements that motivated the project and stated its core contributions. From there, some participants listed subgoals and ideas they could explore (all participants).
In sum, participants used project documents informally to leave a trace of their reasoning and high-level goals, conduct short-term and long-term planning, and progress tracking of these plans.
3.3.2. [RQ2.1] Participants expressed planning intentions in documents as questions or flags
Our analysis of these research documents showed that intention to plan were often expressed using questions or flags (e.g., “TODO” or “[will try]”), and that goals were often related to information seeking (e.g., literature search and review) and other actions (e.g., running experiments). Specifically, many asked themselves a question (often in a how? and can we? format) before creating a plan to answer it. For example, P7 asked “How to construct the corpus?” before creating a plan to detail one approach. While sometimes these questions were written down as a way to provoke thoughts without intent for planning to answer them, flags signaled clear intentions to create a plan. A few participants (P3, P4, P7) also initiated planning from hypotheses they hoped to verify. This can be accomplished through information seeking, running experiments, or both. For example, P3 hypothesized that “similarity matching scores between [texts] describing [concepts] will be higher for [criteria] than those different than [criteria]” and then laid out an experimental plan to verify that hypothesis.
In sum, in all research project documents, we saw places where participants intended to initiate a multi-step plan to address different research tasks. The most common category of tasks are ones that were literature-augmented (e.g., literature search and understanding), which are often combined with a wide variety of other research actions, such as experiment/artifact design (P3), experiment/code execution (P4), data inspection and synthesis (P7), communicating and discussing results (P5), and ideation (P2).
3.3.3. [RQ2.2 & RQ3] Literature-augmented tasks present valuable opportunities for incorporating AI support.
When asked about incorporating AI support into their research documents, all participants expressed a desire to use AI to help with literature-augmented tasks—exploring and understanding relevant literature to inform decision-making. This finding echo recent surveys on how researchers leverage LLM to conduct research, where the most frequent usage category was information-seeking (Liao et al., 2024).
Using the design probe, participants conducted a wide range of high-level to specific literature-grounded tasks, from “what are related works in [field] or [field] about [problem]?” (P9), to“can you list some recent (¿2020) papers that address [problem] for [technology]?” (P2), to “are there any datasets of [domain] or any other interactions between [group] and [group]?” (P4), to “see definitions [of term] from prior papers” (P7).
Participants saw opportunities in leveraging AI systems to save them both time and manual effort for these tasks and to become more effective in these tasks. They pointed to modern AI systems’ ability to retrieve vast documents and to reason over them. P3 explains: “[the WoZ probe] reduces my workload on going and searching on Google Scholar and speeds up [my workflow] more.”, and noted that current search engines do not respond well to natural language questions, which are often how participants planned out literature-augmented tasks in their documents (Section 3.3.2). Further , P4, P5, and P7 mentioned that AI agents today can reason over documents and generate interesting inspirations. Specifically, P9 shared that their goal for a literature search was clarified upon reading AI responses in the probe: “seeing this I’m realizing that what I really wanted to find was a characterization of [concept].”
Overall, participants saw AI as a means of exploring an expansive information space built from literature—perhaps too effortful for them to effectively navigate on their own—and returning resources to inspire new ideas and conclusions.
3.3.4. [RQ3] Participants preferred to perform higher-level reasoning and synthesis themselves
Participants also pointed to tasks in their plans that they did not want AI to automate away. In particular, participants saw higher-level reasoning and information synthesis as critical tasks they wanted to do themselves. They were also often unsatisfied with AI’s outputs on these tasks. P3 explains: “this tinkering process around reading and playing around with things is what gives you the ideas. I don’t know if I want those things automated because the process is as helpful as the final result.” P5 agreed that they “wouldn’t want it to be making the final decisions for sure. Just give me inspiration for where I can go.” Specifically, they shared that “the main thing I worry about is feeling enough ownership if AI makes more consequential decisions. An overeager AI assistant that attempts to perform tasks researchers prefer to do themselves is frustrating because it does not complement the user’s work and instead creates undesired noise for them to filter through.
3.3.5. [RQ3] Document editors are promising environments for human-agent interaction.
@@
Participants also shared how they would like to receive AI assistance. P1 and P2 envisioned an AI system “actively engaging with the content I’m writing […] after I write each statement, the system could retrieve relevant papers that could provide background or related work for me to read more.” Although desirable, participants also identified a challenge with this kind of interaction: context specification. P2 was concerned that they would need to “prompt [the AI] with a lot of background,” but realized that within their project document, they “might have [the background] written down so it’s fine.” For planning in particular, P9 appreciated that during the study, either a facilitator or a participant wrote out plan steps in the document: “the ability to layout the steps when I start something was very helpful [for] visualizing the outcome.” Thus, there is strong motivation for designing an agent to work within and draw from a project document.
Commenting on the separation of the WoZ probe and the planning document and the AI output document, P3 wished for closer integration but still maintaining agency to move output into the user’s workspace when desired: “having something on the side that does not infill into where I am writing, like a side [panel]—if I want something from it and I want to bring it back into my doc, having that agency to decide would be better for the document itself and my experience.” P9 suggested an interaction similar to “in-line comments for Google Docs” to only expose relevant outputs in a sidebar when the information is requested.
In sum, participants wanted tighter integration between an AI agent and their project documents for two main reasons. Firstly, they wanted to receive in-situ AI support as they plan and execute research tasks in their documents, and, secondly, they wanted the AI agent to have access to the broader context that already existed in their project documents.
@@ 3.4. Design Goals
    Integrate seamlessly into a document environment. We learned from answering RQ1 that project documents contain rich externalizations of reasoning. This is key information an AI agent can use to better assist the researcher. We also learned from answering RQ3 that the document is an appealing environment for researcher-agent collaboration, but requires careful information management strategies to prevent unwanted distractions. Thus, we strive to seamlessly embed agents into documents. This involves 1) allowing the researcher to use already-familiar document editing affordances and representations to interact with the AI agent, and 2) strategically managing outputs to avoid excessively cluttering the document.
@@
    Allow flexible delegation of agency between researcher and AI. As we learned from RQs 2 and 3, researchers may not want to delegate all parts of a plan to AI. This preference may also be context-dependent and constantly shifting. Satyanarayan’s calls for flexibly delegating agency between humans and AI (Satyanarayan, 2024) may be one promising approach. We offer a concrete implementation of this flexible delegation in our system.
@@
    Provide opportunities for researcher-AI collaboration in both planning and execution. Delegation is an important aspect of planning, but it does not guarantee successful execution of a plan. We saw throughout the probe activity that researchers encountered (and sometimes even expected) imperfect AI outputs when executing tasks with AI. They then envisioned stepping in to provide the agent with more context or iterating on the outputs themselves. We thus see two stages for fertile researcher-AI collaboration: planning (collaboratively devising a plan to tackle a problem) and execution (collaboratively completing steps in the plan). In practice, these two stages are closely intertwined and synergistic, so we aim to support both simultaneously.
@@ 4. Cocoa: System Walkthrough and Implementation
In light of our design goals, we present Cocoa, a system that embeds an AI agent into a document editor using a novel interaction design pattern we call interactive plans. Interactive plans allow users to collaboratively plan (co-plan) with the agent—the agent proposes an initial plan of execution to tackle a user request that the user can edit to their liking. Then, users can collaboratively execute (co-execute) the plan with the agent—the user and the agent can build off each others’ work to synergize human and AI capabilities, and provide flexible negotiation of human and machine agency.
@@ 4.1.1. Invoking the agent and selecting plans
Nóírín’s project document is an informal, reverse chronological log of research progress and includes information such as meeting notes, rough ideas, links to literature, and questions for herself and her collaborators. To initiate co-planning, Nóírín can highlight any piece of text in her document that (implicitly or explicitly) contains a request she would like to tackle with the agent. She highlights a question she previously wrote but has not gotten an opportunity to explore yet: “What are new ways AI agents can interactively elicit human feedback?” She then invokes the agent with a button that appears above the highlighted text in a floating menu. The agent analyzes the request within the context of other text in her document and proposes a series of plans for answering the question. These plans are displayed in the document via a plan selector UI within the editor, which allows Nóírín to browse and select a plan. Upon Nóírín’s selection, a plan becomes fully interactive within the document. This interactive plan is comprised of a series of editable plan steps as a bulleted list. Each plan step further consists of several interactive components.
@@ 4.1.2. Agent steps and user steps
Nóírín can use the step assignment toggle to assign the step to herself (a “user step”) or the agent (an “agent step”). User steps are highlighted in blue. When executing the plan, the agent will automatically attempt agent-assigned steps but will request Nóírín’s input on user-assigned steps; when that happens, a step will be highlighted in orange. For now, Nóírín is satisfied with how the agent has left some steps that require higher level reasoning to herself.
@@ 4.1.3. Editing the plan step description
The step description appears in an editable text field and describes a subtask in natural language while doubling as high-level instructions for the agent. Nóírín can edit the step description just like editing any other bulleted list in her document. The system will prompt Nóírín to save any changes upon editing with the keyboard shortcut Cmd/Ctrl+S. In addition to editing the step description herself, Nóírín can also employ LLM assistance—highlighting any text within a step description will present a button in a floating menu labeled “Suggest alternate step.” When clicked, an LLM suggests a step to replace the current one, informed by the previous steps. Nóírín tries this a few times on a step she wishes to modify, and although the agent’s suggestions are interesting, she feels like the original step is still closest to what she prefers. Nóírín can easily undo these suggestions with the keyboard shortcut Cmd/Ctrl+Z.

### v3

@@ 2.3. AI for Scientific Research
As AI systems become more capable, we ask this question: Is the generation of full research artifacts (e.g., research questions, proposals) the most desired and promising paradigm to assist researchers? Indeed, when working with other (human) collaborators, researchers often find richness in co-evolving partially completed artifacts (Star and Griesemer, 1989; Lee, 2007; Shrum et al., 2007; Jirotka et al., 2013) and answering or asking questions that stimulate critical thinking and reflection (Overholser, 1993; Padesky, 1993). Further, some fully automatically generated artifacts are still found to be of lower quality than those generated with human involvement (Jansen et al., 2025). A limited number of works have thus explored human-AI co-creation of research artifacts (Babaian et al., 2002; Liu et al., 2024a; Nigam et al., 2024; Kang et al., 2022b). These include in-document commands that trigger an AI assistant to complete a partial citation based on the user’s preferred bibliographies and paper collections (Babaian et al., 2002), interactive node-based editors to iteratively expand upon and refine AI-generated research questions (Liu et al., 2024a; Pu et al., 2024), and emulating colleague and mentor personas using LLMs to work with the user in developing research proposals (Nigam et al., 2024).
@@
However, despite these systems, there is little clarity on how researchers want to co-create with AI during research planning and execution. Prior work (e.g., (Liu et al., 2024a; Nigam et al., 2024)) assumed that generating particular artifacts is helpful to researchers. Rather than focusing on generating predetermined research artifacts, we contribute an interaction design pattern that allows for flexible specification of the final artifact and collaboration between a researcher and an AI agent.
@@ 3. Formative Study
Building on prior work, we aim to enhance collaboration between humans and AI agents using a shared, interactive operational representation. We focus on scientific research—a challenging domain for planning given its complex, multi-step workflows and the nuanced decisions researchers make upon processing information-dense scientific papers. Specifically, we investigate the potential for project documents—continuous records of ideas, updates, and tasks—to act as the agent environment. We thus conducted a formative study to answer the following research questions:
  • F1:
    What are the properties and opportunities of researchers’ project documents for human-agent collaboration?
  • F2:
@@
  • F3:
    How would researchers prefer to collaborate with an AI agent in their research workflows?
3.1. Participants, Procedure, and Data Analysis
We recruited 9 Ph.D. students (detailed demographics in Table 1 of Appendix B) through an interest form sent to a research organization’s internal Slack channel and the authors’ professional connections. We targeted Ph.D. students as they often lead the detailed planning and execution of research projects and are the primary editors of project documents.
All studies were conducted virtually over Google Meet and lasted 60 minutes. Before the study, we collected from participants an active or past research project document and a brief description of their research interests. The study was divided into 3 parts:
  1. (1)
    An activity involving the participant’s project document to better understand their current planning behavior (25 minutes).
  2. (2)
    An activity with a Wizard-of-Oz (WoZ) design probe to explore how researchers plan and explore research ideas in a document editor with LLM support via a chatbot (25 minutes).
  3. (3)
    An exit interview where researchers reflected on their experience with the probe and using LLMs in research (10 minutes).
Details for each part can be found in Appendix C. Each participant was given a \$35 USD honorarium after the study. The study was reviewed and exempted by our organization’s internal review board. We qualitatively analyzed project documents produced from the study probe activity alongside study transcripts. Our data analysis methods can be found in Appendix D.
@@ 3.2. Findings
In this section, we redact any project-specific details for privacy and intellectual confidentiality.
3.2.1. [F1 & F3] Project documents are promising environments for human-agent interaction
Participants’ project documents acted as a “hub” for their projects. These documents commonly included meeting notes, planned to-do items, progress updates tied to to-do items, research questions to discuss with collaborators, and links to other documents detailing a particular project component in greater depth. These documents often served as “scratch paper” for researchers to informally leave a trace of their reasoning and high-level goals, conduct short-term and long-term planning, and progress tracking of these plans. Most documents (P1, P2–4, P7–P9) contained one or more problem statements that motivated the project and stated its core contributions. From there, participants listed subgoals and ideas for planned exploration (all participants).
@@
Participants expressed a desire for tighter integration between an AI agent and their project documents in their research workflows for two main reasons. First, they wanted to receive in-situ AI support as they plan and execute research tasks in their documents. P1 and P2 both envisioned an AI system “actively engaging with the content I’m writing […] after I write each statement, the system could retrieve relevant papers that could provide background or related work for me to read more.” P9 appreciated “the ability to lay out the steps when I start something. [It] was very helpful [for] visualizing the outcome.” Second, participants wanted the AI agent to have access to the broader context that already existed in their project documents. When using the probe, P2 was concerned that they would need to “prompt [the chatbot] with a lot of background,” but realized they “might have [the background] written down” in their document. P3 desired closer document integration but warned that the agent may clutter their document. They suggested: “having something on the side that does not flow into where I am writing, like a side [panel]—if I want something from it, I want to bring it back into my doc.”
3.2.2. [F2] Literature search and understanding play a central role in research planning
In all project documents, we saw places where participants intended to initiate a multi-step plan (often via a how? and can we? question) to address different research tasks. The most common category of tasks are ones that were literature-augmented (e.g., literature search and understanding), which are often combined with a wide variety of other research actions, such as experiment/artifact design (P3), experiment/code execution (P4), data inspection and synthesis (P7), communicating and discussing results (P5), and ideation (P2). Literature was not only a significant component of existing plans, but was important for informing future plans. Some participants’ plans (e.g., P3, P4, P7) were formed from hypotheses they hoped to verify, which prior work may have already addressed; P7 pointed out that the hypotheses they had in mind can “actually be proven from previous papers” so they saw literature review as a planning aid. Consequently, all participants expressed a desire to use AI to help with literature-augmented tasks—exploring and understanding relevant literature to inform decision-making. This finding echoes recent surveys on how researchers leverage LLMs to conduct research, where the most frequent usage category was information-seeking (Liao et al., 2024).
3.2.3. [F3] Participants preferred to perform higher-level reasoning and synthesis themselves
In addition to highlighting tasks where they wanted AI assistance, participants also spoke about tasks they did not want AI to automate away. In particular, participants saw higher-level reasoning and information synthesis as critical tasks they wanted to do themselves. They were also often unsatisfied with AI’s outputs on these tasks. P3 explains: “this tinkering process around reading and playing around with things is what gives you the ideas. I don’t know if I want those things automated because the process is as helpful as the final result.” P5 agreed that they “wouldn’t want it to be making the final decisions for sure. Just give me inspiration for where I can go.” Specifically, they shared that “the main thing I worry about is feeling enough ownership” if AI makes more consequential decisions. An overeager AI assistant that attempts to perform tasks researchers prefer to do themselves is frustrating because it does not complement the user’s work and instead creates undesired noise for them to filter through.
@@ 3.3. Design Goals
    Integrate seamlessly into a document environment. Project documents contain rich externalizations of researchers’ workflow, thought processes, and plans (e.g., “to-dos”) [F1]. This is key information an AI agent can use to better assist the researcher. The document is also an appealing environment for researcher-agent collaboration, but requires careful information management strategies to prevent unwanted distractions [F3]. Thus, we aim to 1) allow the researcher to use familiar document editing affordances to interact with the agent, and 2) strategically manage outputs to avoid excessively cluttering the document.
@@
    Allow flexible delegation of agency between researcher and agent. Researchers may not want to delegate all parts of a plan to AI [F2, F3]. This preference may also be context-dependent and constantly shifting. Our system uses a concrete implementation of a flexible delegation approach proposed by Satyanarayan (Satyanarayan, 2024).
@@
    Provide opportunities for researcher-AI collaboration in both planning and execution. Agent developers have called for decoupling planning and execution to avoid fruitless execution (Huyen, 2025). Indeed, participants discussed planning (i.e., breaking down a problem into smaller, more actionable tasks like they did in their project documents) and execution (i.e., completing steps in the plan and refining outputs, with or without AI assistance) as separate but complementary actions. We thus see planning and execution as two distinct stages for fertile researcher-AI collaboration. Because these two stages are closely intertwined and synergistic, we aim to interleave the two in our system.
@@ 4. Cocoa: System Walkthrough and Implementation
We present Cocoa, a system that embeds an AI agent into a document editor using a novel interaction design pattern we call interactive plans. Interactive plans allow users to collaboratively plan (co-plan) with the agent—the agent proposes an initial plan of execution to tackle a user request that the user can edit to their liking. Then, users can collaboratively execute (co-execute) the plan with the agent—the user and the agent can build off each other’s work to synergize human and AI capabilities, and provide flexible negotiation of human and machine agency.
@@ 4.1.1. Invoking the agent and selecting plans
Nóírín’s project document is an informal, reverse chronological log of research progress and includes information such as meeting notes, rough ideas, links to literature, and questions for herself and her collaborators. She has a specific question in her document that she wants to explore further, but is not sure how to proceed on her own: “What are new ways AI agents can interactively elicit human feedback?”
To initiate co-planning, she highlights the question to invoke the agent on it using a button that appears in a floating menu. The agent analyzes the question within the context of her entire document and proposes a few different plans for approaching it. These plans are displayed in the document via a plan selector UI within the editor, which allows Nóírín to browse and select a plan. Upon Nóírín’s selection, a plan becomes fully interactive within the document (Figure 3).
@@
[…]
A user invokes the agent on a piece of text in the document by clicking on the “Invoke agent” button that appears on hover whenever text is highlighted. The agent will use the highlighted text and context from elsewhere in the document to propose a series of plans, displayed in a plan selector that appears under the highlighted text. Once the user selects a plan, it becomes fully interactive in the document.
@@ 4.1.2. Agent steps and user steps
The initial plan included steps for brainstorming relevant topics, searching for papers, and making connections between papers (e.g., Figure 3). Nóírín wants to assign tasks that are relatively low risk but high effort to the agent, such as searching for papers and expanding on some of her preliminary ideas, and keep tasks requiring higher-level thinking and deep research expertise to herself, such as identifying papers that can seed deeper exploration and surfacing connections between papers. To do so, she uses the step assignment toggle (Figure 4) to assign the step to herself (a “user step”) or the agent (an “agent step”). User steps are highlighted in blue. When executing the plan, the agent will automatically attempt agent steps but will request Nóírín’s input on user steps.
@@
[…]
An agent step (plan step assigned to the agent) is shown above a user step (plan step assigned to the user). The user can assign the step to either themselves or the agent by using the step assignment toggle beside the step, as well as edit the step’s description directly, just like they were editing text elsewhere in the document.
@@ 4.1.3. Editing the plan step description
Upon further review of the plan, Nóírín is now curious about authors who have published on “interactive feedback mechanisms in AI” rather than seeing papers on the topic directly. She edits the step description to reflect this. The step description can be edited just like editing any other bulleted list in her document, making this intuitive for Nóírín. If she chooses, Nóírín can also employ LLM assistance to suggest a step to replace the current one, informed by the previous steps (Figure 5). Nóírín can easily undo the suggestion if she prefers the original step. Once Nóírín is satisfied with her editing, she saves her changes with the keyboard shortcut Cmd/Ctrl+S to register them with Cocoa.
@@
[…]
The user highlights text within a plan step, revealing a button labelled ’Suggest alternate step.’ Clicking the button would replace the current plan step description with an AI-generated suggestion.